is 6+4x=3y parallel, perpendicular or neither to 3x-4y=8?
Answer by Flannery(124)

6+4x=3y;3x-4y=8
Find slope intercept form of those equations.
(Meaning solve them for y.)
y=4/3x+2;y=3/4x-2
Point slope form is y=bx+c where x and y are variables and where b and c are constants.
If the b terms in two equations are the same the lines are parallel. If the b terms are negative reciprocals the lines are perpendicular.
The b terms here are 3/4 and 4/3. Reciprocals but not negative reciprocals.
So the answer is neither.
Checked with Microsoft Mathematics.